40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 17 07:35:48 2022, yuta, Bureaucracy, General, My wish list for IFO commissioning 
    Reply  Fri Aug 19 14:46:32 2022, Anchal, Update, SUS, Open loop transfer function measurements for local damping loops of BHD optics 
       Reply  Sat Aug 20 20:26:10 2022, Anchal, Update, SUS, Open loop transfer function measurements for local damping loops of Core optics ALL_CORE_SUS_OLTF_2022-08-20-merged.pdf
          Reply  Mon Aug 22 14:36:49 2022, rana, Update, SUS, Open loop transfer function measurements for local damping loops of Core optics 
Message ID: 17085     Entry time: Wed Aug 17 07:35:48 2022     Reply to this: 17092
Author: yuta 
Type: Bureaucracy 
Category: General 
Subject: My wish list for IFO commissioning 

FPMI related
- Better suspension damping HIGH
 - Investigate ITMX input matrix diagonalization (40m/16931)
 - Output matrix diagonalization
 * FPMI lock is not stable, only lasts a few minutes for so. MICH fringe is too fast; 5-10 fringes/sec in the evening.
- Noise budget HIGH
 - Calibrate error signals (actually already done with sensing matrix measurement 40m/17069)
 - Make a sensitivity curve using error and feedback signals (actuator calibration 40m/16978)
 * See if optical gain and actuation efficiency makes sense. REFL55 error signal amplitude is sensitive to cable connections.
- FPMI locking
 - Use CARM/DARM filters, not XARM/YARM filters
 - Remove FM4 belly
 - Automate lock acquisition procedure
- Initial alignment scheme
 - Investigate which suspension drifts much
 - Scheme compatible with BHD alignment
 * These days, we have to align almost from scratch every morning. Empirically, TT2 seems to recover LO alignment and PR2/3 seems to recover Yarm alignment (40m/17056). Xarm seems to be stable.
- ALS
 - Install alignment PZTs for Yarm
 - Restore ALS CARM and DARM
 * Green seems to be useful also for initial alignment of IR to see if arms drifted or not (40m/17056).
- ASS
 - Suspension output matrix diagonalization to minimize pitch-yaw coupling (current output matrix is pitch-yaw coupled 40m/16915)
 - Balance ITM and ETM actuation first so that ASS loops will be understandable (40m/17014)
- Suspension calibrations
 - Calibrate oplevs
 - Calibrate SUSPOS/PIT/YAW/SIDE signals (40m/16898)
 * We need better understanding of suspension motions. Also good for A2L noise budgeting.
- CARM servo with Common Mode Board
 - Do it with single arm first

BHD related
- Better suspension damping HIGH
 - Invesitage LO2 input matrix diagonalization (40m/16931)
 - Output matrix diagonalization (almost all new suspensions 40m/17073)
 * BHD fringe speed is too fast (~100 fringes/sec?), LO phase locking saturates (40m/17037).
- LO phase locking
 - With better suspensions
 - Measure open loop transfer function
 - Try dither lock with dithering LO or AS with MICH offset (single modulation)
 - Modify c1hpc/c1lsc so that it can modulate BS and do double demodulation, and try double demodulation
- Noise Budget HIGH
 - Calibrate MICH error signal and AS-LO fringe
 - Calibrate LO1, LO2, AS1, AS4 actuation using ITM single bounce - LO fringe
 - Check BHD DCPD signal chain (DCPD making negative output when fringes are too fast; 40m/17067)
 - Make a sensitivity curve using error and feedback signals
- AS-LO mode-matching 
 - Model what could be causing funny LO shape
 - Model if having low mode-matching is bad or not
 * Measured mode-matching of 56% sounds too low to explain with errors in mode-matching telescope (40m/16859, 40m/17067).

IMC related
- WFS loops too fast (40m/17061)
- Noise Budget
- Investigate MC3 damping (40m/17073)
- MC2 length control path

ELOG V3.1.3-